# From reductionism to realism: Holistic mathematical modelling for complex biological systems

### v1

@@ Abstract
At its core, the physics paradigm adopts a reductionist approach to modelling, aiming to understand fundamental phenomena by decomposing them into simpler, elementary processes. While this strategy has been tremendously successful in physics and is typically considered the pinnacle of scientific formulation, it has often fallen short in addressing fundamental questions in the biological sciences. This limitation arises from the inherent complexity of biological systems, characterised by heterogeneity, poly-functionality and interactions across multiple spatial and temporal scales. Nevertheless, the traditional framework of complex systems modelling also falls short, as its emphasis on broad theoretical principles has often failed to produce realistic, predictive, empirically grounded insights. To advance towards the goal of actionable mathematical models in biology, we argue here, using neuroscience as a case study, that it is necessary to move beyond simple reductionist approaches and instead embrace the intrinsic complexity and heterogeneity of biological systems—leveraging the growing availability of high-resolution data and recent advances in high-performance computing. In particular, we advocate for a holistic mathematical modelling paradigm that harnesses rich representational structures such as annotated and multilayer networks, employs agent-based models and simulation-based approaches, and focuses on the inverse problem of inferring system properties from dynamical observations. Finally, we emphasise that this approach is fully compatible with the search for fundamental principles in biophysics, and we highlight the substantial potential it holds to drive progress in mathematical biology over the next two decades.
@@ I Introduction
A fundamental aspect of human nature and success as a species is the desire to make sense of the world around us. Over the course of thousands of years and through the scientific revolution, mathematics, specifically mathematical modelling (MM), has become the dominant and ultimate paradigm used to abstract physical systems and describe their behaviour over time and space. This has resulted in the unparalleled success of MM in domains like physics and engineering, and has profoundly changed human society. The relationship has become naturally symbiotic with new mathematics constantly being discovered through the development of models of real-world phenomena. Many such examples have become canon in the field of MM such as Poincaré’s attempts to model celestial mechanics [1], which laid the ground-work for the study of dynamical systems, divergent series, and resonance phenomena. In an effort to understand the bridges of Koenigsberg, Euler developed concepts that would later culminate in graph and network theory [2]. More recently, whilst deriving simple equations to describe the dynamics of the climate, Lorenz discovered the first example of a chaotic attractor [3]. Whilst we are still far from a complete theory, it can be argued ${}^{\text{th}}$ century was the century of physics when mathematical abstractions led to a significant maturity in the field, with many fundamental problems being solved and theories confirmed [4].
@@
Nevertheless, it would be inaccurate to characterize mathematical biology as an immature discipline, as numerous models have demonstrated great success and provided valuable insights. However, these breakthroughs often exist in isolation and are typically limited to specific biological systems or processes. To understand why this limitation arises from the intrinsic nature of biological systems, we must first identify the characteristics of the dominant reductionist approach that has resulted in such rampant success in physics. Throughout this Perspective, we will built our arguments with examples primarily from neuroscience and the human brain, which serves as the prototypical example of a complex biological system due to its inherent complexity, intricate organization, and central importance to all aspects of human life and society.
@@
The reductionist approach has proved extremely successful across science, but particularly in physics and chemistry, where simple experiments with only a few independent variables can reveal fundamental principles, and fundamental principles can be assembled together like building blocks to explain more complex situations. However, this approach may also be responsible for the lack of theoretical success that has plagued the biological sciences [7, 8, 9, 10]. For example, whilst there are only a few elementary units for physical systems e.g. protons, neutrons and electrons, biological systems are characterised by an array of unique elements e.g. the zoo of different lymphocytes in the immune system, the collections of proteins, or the extremely diverse populations of neurons and glial cells in the nervous system. Returning to Hodgkin and Huxley, whilst their result represented the crucial breakthrough in understanding the subcellular mechanisms governing neural firing, it brought us only marginally closer to understanding cognition or computational processes in the brain.
@@
In this Perspective, we focus on the problem of biological complexity and MM. In particular we argue that the reductionist approach must be avoided in order to build a holistic approach that develop models consolidating different spatial and temporal scales, using multimodal data, and considering complex representational structures. This holistic mathematical modelling philosophy goes beyond isolated functions to integrate multiple biophysical processes within a single model. Moreover, we make the case that the theoretical approaches to complex systems that have been dominant for the past two decades, are insufficient to yield empirically useful or actionable insights. Looking to the future, we argue that modern tools in data collection, machine learning (ML), artificial intelligence (AI), and high-performance computing (HPC) will not replace modelling, but instead facilitate it at unprecedented scale. By adopting a modern approach, we make the case for a holistic modelling framework that embraces complexity and empiricism in order to answer the most pressing scientific challenges.
@@ II Rich and realistic representations for complex systems
In an effort to accurately represent real-world systems, from biology and beyond, the study of networks has become one of the most popular methods for abstracting elements and their interactions [18]. In the last 20 years, this has lead to a number of important studies such as that of ‘scale-free’ [19], ‘small-world’ [11], or community-structured networks [20]. As a result, the study of complex systems has become, for many, synonymous with network science. This abstraction has been particular useful in the modelling of a range of biological systems such as the brain [21], ecosystems [22], biomedical systems [23], protein interactions [24] and gene regulation [25]. By introducing weighted, directed and signed connections, as well as temporal networks, where links vary in time, these abstractions can be extended to model a variety of different biological processes.
@@
Whilst intrinsically motivated by applications, the study of complex systems has, to date, remained largely theoretical, with a substantial focus on general principles that aim to transcend the confines of a specific system or process [26]. More recently, significant attention has turned towards higher-order representations of complex systems [27, 28, 29], which generalise networks to more expressive structures such as hypergraphs, simplicial complexes and multilayer networks, illustrated in Fig. 2. In particular, multilayer networks present new opportunities for modelling biological systems where a number of different, yet coupled processes, are evolving simultaneously [30]. For example, neural activity in the human brain network is facilitated by blood-flow through the cerebral vasculature. Moreover, ageing results in structural changes to the networks that faciliate cognition. Whilst these processes have different mechanisms, dynamics and time-scales, they remain intrinsically coupled, reducing the efficacy of any model that attempts to consider them in isolation.
@@
Somewhat surprisingly, one higher-order representation of a complex system whose dynamics are rarely considered is the annotated network [^1] It is worth noting that whilst annotated networks have been rarely studied in terms of dynamics, they are prevalent in some fields, such as graph-representation learning [31]. [32]. Unlike the simple structure found in graph theory of a typical network, $G=(V,E)$, defined by a vertex set, $V$, and edge set, $E$, annotated networks can encode nodal metadata in the form of an additional vector of features $\mathbf{d}_{i}\in\mathbb{R}^{d}$ that is attached to each vertex $i\in V$ (or to each edge) [33, 34]. In particular, this extension allows us to model the heterogeneity of individual elements, adding a crucial layer of additional expressiveness to our mathematical models. Whilst still nascent, such an approach has proved useful in neuroscience in the form of biologically annotated connectomes [35], which are human brain networks obtained from structural imaging where each node (region) is imbued with additional information such as cell-type density, gene transcription, receptor density, and myelination. Open-source software, such as neuromaps [36], facilitates the integration of this multimodal data into richer structural representations. Such heterogeneity has been used to develop biophysical models of whole-brain dynamics that incorporate regional heterogeneity through variations in gene expression [37] or synaptic strength [38] for added biological complexity and closer agreement with observed neural dynamics.
@@
In an attempt to tackle the complexity of biological systems, the time has come to apply these frameworks to concrete problems in biological domains. In turn, such fields may replace reductionist abstractions with richer representations that are capable of incorporating the heterogeneity seen in experiments. In the past, such over-parametrised models were plagued by both a lack of sufficient multimodal data and the inference algorithms needed to fit them. However, as we progress into the age of big data, modern experimental techniques are able to record multimodal data on an unprecedented scale, particularly in biological systems [39, 40]. Moreover, novel techniques in computational statistics and ML are able to perform parameter inference for such complex models using observed data [41, 42]. With accurate calibration on experimental data and efficient simulation on HPCs, detailed, heterogeneous and elaborate models of complex systems can make the leap from descriptive to mechanistic and predictive.
@@ II-A Integrating many time- and length-scales
The situation is completely different for biological systems that are best characterised by interacting spatiotemporal scales [43]. Focusing on a single scale, naturally limits our ability to explain certain phenomenon. For example, neurodegenerative diseases, like Alzheimer’s disease (AD), evolve due to the spread of toxic proteins over the course of many years. However, the symptoms of these conditions manifest themselves in cognitive decline stemming from pathological neural activity, on the order of $10^{-3}-10^{0}$ seconds [44]. In order to understand the mechanisms by which AD causes cognitive decline, it is insufficient to only model either neural activity or toxic protein spread in isolation. Fig. 3 highlights the many processes underlying the development of AD along with their respective time- and length-scales. Recent studies have presented an approach that attempts to integrate these time-scales [45, 46]. First, they model network degradation due to toxic-protein spread, freezing the state of the network at distinct snapshots to then form the topology underlying a model of fast neural activity.
@@
Figure 3: Spatiotemporal scales in Alzheimer’s disease. Complex biological processes, such as the development of neurodegenerative diseases such as Alzheimer’s disease (AD), evolve over a range of interacting spatiotemporal scales. Focussing on AD, protein misfolding ultimately leads to the spread of toxic tau and amyloid-$\beta$ proteins through axonal fibres, which is partially cleared through cell regulation. Ultimately, this leads to cell death and pathological neural dynamics. This manifests in tau and amyloid pathology and, finally, cognitive decline which can be combatted with therapy. A MM of this process must take into account the full range of spatiotemporal scales in order to build a mechanistic understanding of AD.
@@ III-A Agent-based models
Mathematical models are often appreciated for their elegance and parsimony. In particular, such models can often be treated analytically leading to great insight. As models become more complex, they lose analytical tractability and numerical methods of analysis must be used. One framework for modelling complex systems that hinges on direct simulation, is the use of agent-based models (ABMs), which have been embraced in economic research [47, 48], with some use in mathematical biology [49, 50].
The typical structure of ABMs is the time evolution of a micro-scale model that defines the characteristics of individual elements and their interactions. Examples of such systems include models of the spread of disease through populations and ecological dynamics or collective motion with agents being individual or populations (Panel A, Fig. 4). Unlike traditional models, ABMs call for a different approach to mathematical analysis, one that is often realised through exhaustive simulation and data-driven calibration. For this reason, such models were previously found to be computationally restrictive. However, following advancements in scientific computing, large-scale simulations of ABMs have become both tractable and attractive [51].
The adoption of ABMs requires a shift in attitudes towards mathematical models. In particular, the proponents of such models argue that the complexity of many systems requires a shift towards a purely simulation-based science, where additional features and realism are favoured over parsimony and elegance [52]. The current inclination toward a minimal model is often justified via Occam’s razor, the philosophical argument that the simplest model is often the best. However, the lack of success of such minimal models in biology, compounded with the success of overparameterised ML models, has called into question this long-standing perspective [53, 54]. Whilst ML models are able to discover a multiplicity of latent features that have predictive power, a parsimonious mathematical modeller must instead identify the handful that may be important. Whilst not without critics [55], more complex models relieve the modeller of some of this choice by allowing them to include sufficient complexity at the level of model definition. Moreover, it is important to note the differences between an ABM and a ML model. Whilst both lack analytical tractability and rely on computationally-intensive processes, the latter may be considered to be a ‘black-box’, whose mechanisms remain somewhat opaque, whilst the former attempts to model the basic physical mechanisms governing a system. This poses additional advantages, such as the ability to test realistic interventions in a system due to the model’s interpretability.
@@ III-B Digital twins
Another simulation-based framework for the modelling of complex systems, is the digital twin (DT). A DT is an adaptive model that is designed to virtually emulate the behaviour of a physical system [58]. Typically such a model is implemented on a computer. DTs have been considerably useful in engineering where they can be used to simulate the behaviour of a system under a range of simulated conditions. Successful examples include spacecraft during the Apollo missions [59], as well as other mechanical and aerospace engineering systems [60]. In some cases, a DT is composed of an ABM, such as in models of cities. Similarly to ABMs, whilst DTs have an underlying mathematical formulation, they require exhaustive simulation and HPC. Moreover, they are often characterised by more engineered interfaces such as virtual or augmented reality (VR/AR), for use by stakeholders other than the mathematical modeller.
@@
In biology, DTs have sparked particular interest in the field of precision medicine [61]. Precision medicine refers to a treatment paradigm that considers the individual characteristics and data of a patient. This can be realised through the use of a personalised DT that uses biomedical data to calibrate a simulation of a specific patient’s system to test interventions in-silico [^2] As opposed to experiments in-vitro (in glass i.e. in a laboratory) or in-vivo (in a living system), in-silico refers to a third paradigm where experiments can be run on a computer., as illustrated in Panel B, Fig. 4. Two particular systems showing great promise are the heart [62] and the brain [63]. In particular, The Virtual Brain (TVB), more specifically The Virtual Epileptic Patient (VEP), is a project leveraging detailed biophysical models and HPCs, that predicts personalised estimates of epileptogenic zone networks [64]. The empirical, actionable success of this approach stands in contrast with the dominant paradigm of complex systems neuroscience that focuses on simple oscillatory models of neural dynamics [65] and has yet to realise a concrete medical application [66]. Nevertheless, it would be naive to assume that the pure emulation of a system is either sufficient or necessary to understand it. The Human Brain Project [^3] Continuing now as the Blue Brain Project., was one such attempt to emulate the whole brain from individual neurons upwards [67], that, despite some notable results at the micro-scale [68], failed to deliver on its original promises [69]. Instead, DTs should incorporate the necessary complexity to answer specific questions, test interventions, and simulate different environments, limited by the available data, rather than a perfectly faithful reconstruction of an entire system which achieves very little in scientific terms.

### v2

@@ Abstract
At its core, the physics paradigm adopts a reductionist approach, aiming to understand fundamental phenomena by decomposing them into simpler, elementary processes. While this strategy has been tremendously successful in physics, it has often fallen short in addressing fundamental questions in the biological sciences. This arises from the inherent complexity of biological systems, characterised by heterogeneity, polyfunctionality and interactions across spatiotemporal scales. Nevertheless, the traditional framework of complex systems modelling falls short, as its emphasis on broad theoretical principles has often failed to produce predictive, empirically-grounded insights. To advance towards actionable mathematical models in biology, we argue, using neuroscience as a case study, that it is necessary to move beyond reductionist approaches and instead embrace the complexity of biological systems—leveraging the growing availability of high-resolution data and advances in high-performance computing. We advocate for a holistic mathematical modelling paradigm that harnesses rich representational structures such as annotated and multilayer networks, employs agent-based models and simulation-based approaches, and focuses on the inverse problem of inferring system dynamics from observations. We emphasise that this approach is fully compatible with the search for fundamental biophysical principles, and highlight the potential it holds to drive progress in mathematical biology over the next two decades.
@@ I Introduction
[…]
@@
Nevertheless, it would be inaccurate to characterise mathematical biology as an immature discipline, as numerous models have demonstrated great success and provided valuable insights. However, these breakthroughs often exist in isolation and are typically limited to specific biological systems or processes. To understand why this limitation arises from the intrinsic nature of biological systems, we must first identify the characteristics of the dominant reductionist approach that has resulted in such rampant success in physics. Throughout this Perspective, we will build our arguments with examples primarily from neuroscience and the human brain, which serves as the prototypical example of a complex biological system due to its inherent complexity, intricate organisation, and central importance to all aspects of human life and society.
@@
The reductionist approach has proved extremely successful across science, but particularly in physics and chemistry, where simple experiments with only a few independent variables can reveal fundamental principles, and fundamental principles can be assembled together like building blocks to explain more complex situations. However, this approach may also be responsible for the lack of theoretical success that has plagued the biological sciences [7, 8, 9, 10, 11]. For example, whilst there are only a few elementary units for physical systems e.g. protons, neutrons and electrons, biological systems are characterised by an array of unique elements e.g. the zoo of different lymphocytes in the immune system, the collections of proteins, or the extremely diverse populations of neurons and glial cells in the nervous system. Returning to Hodgkin and Huxley, whilst their result represented the crucial breakthrough in understanding the subcellular mechanisms governing neural firing, it brought us only marginally closer to understanding cognition or computational processes in the brain.
@@
In this Perspective, we focus on the problem of biological complexity and MM. In particular, we present a unified framework of complementary research directions, which together form what we call holistic mathematical modelling. We make the case that in modern sciences the classical reductionist approach cannot work in isolation, and it is necessary to develop models that consolidate different spatial and temporal scales, use multimodal data, and consider complex representational structures. This holistic approach goes beyond isolated functions to integrate multiple biophysical processes within a single model. Moreover, we argue that the theoretical approaches to complex systems that have been dominant for the past two decades, especially those models that are inspired by biological problems, are limited and often cannot yield empirically useful or actionable insights. Yet, the extent of the data and computational requirements needed for high-fidelity models is often prohibitive. Despite this, we believe that the development of such models is a necessity for progress in mathematical biology to match the accomplishments of experimental biology. Looking to the future, we argue that modern tools in data collection, machine learning (ML), artificial intelligence (AI), and high-performance computing (HPC) will not replace modelling, but instead facilitate it at unprecedented scale. By adopting a modern approach, we present a holistic modelling framework that embraces complexity and empiricism, whilst synergising with exisiting approaches, to develop a collective arsenal capable of answering the most pressing scientific challenges.
@@ II Rich and realistic representations for complex systems
In an effort to accurately represent real-world systems, from biology and beyond, the study of networks has become one of the most popular methods for abstracting elements and their interactions [19]. In the last 20 years, this has lead to a number of important studies such as that of ‘scale-free’ [20], ‘small-world’ [12], or community-structured networks [21]. As a result, the study of complex systems has become, for many, synonymous with network science. This abstraction has been particularly useful in the modelling of a range of biological systems such as the brain [22], ecosystems [23], biomedical systems [24], protein interactions [25] and gene regulation [26]. By introducing weighted, directed and signed connections, as well as temporal networks, where links vary in time, these abstractions can be extended to model a variety of different biological processes.
Whilst intrinsically motivated by applications, the study of complex systems has, to date, remained largely theoretical, with a substantial focus on general principles that aim to transcend the confines of a specific system or process [27]. Similarly, many studies aimed directly at biological complexity, focus on ‘top-down’ models, which aim at reproducing qualitative emergent phenomena [28], as opposed to predictive accuracy. Whilst this is no doubt important, it has broadly failed to represent biological systems with the required authenticity to posit or confirm theories, or often to yield practical results. As a result, complex systems science can often exist in a vacuum, adjacent to the mainstream of biology.
In recent times, in an effort to generalise networks to more expressive, realistic structures,
@@
significant attention has turned towards higher-order representations of complex systems [29, 30, 31] such as hypergraphs, simplicial complexes and multilayer networks, illustrated in Fig. 2. Multilayer networks have proved to be particularly useful representations for multi-scale biological systems [32] as they allow for modelling a number of different, yet coupled processes, evolving simultaneously [33, 34]. For example, neural activity in the human brain network is facilitated by blood-flow through the cerebral vasculature. Moreover, ageing results in structural changes to the networks that facilitate cognition. Whilst these processes have different mechanisms, dynamics and time-scales, they remain intrinsically coupled, reducing the efficacy of any model that attempts to consider them in isolation. Despite this, many models on multilayer networks remain largely theoretical, as they are typically overparameterised, requiring large amount of multimodal data for fitting. Nevertheless, the increasing availability and quality of multimodal data raises the opportunity for this expressive modelling technique to be integrated into more powerful, predictive methods, as shown by the recent MINT toolbox for neuroimaging data in disease [35]. We believe that the next 20 years will see an increased focus on this type of data and modelling.
@@
Somewhat surprisingly, one higher-order representation of a complex system whose dynamics are rarely considered is the annotated network [^1] It is worth noting that whilst annotated networks have been rarely studied in terms of dynamics, they are prevalent in some fields, such as graph-representation learning [36]. [37]. Unlike the simple structure found in graph theory of a tyical network, $G=(V,E)$, defined by a vertex set, $V$, and edge set, $E$, annotated networks can encode nodal metadata in the form of an additional vector of features $\mathbf{d}_{i}\in\mathbb{R}^{d}$ that is attached to each vertex $i\in V$ (or to each edge) [38, 39]. In particular, this extension allows us to model the heterogeneity of individual elements, adding a crucial layer of additional expressiveness to our mathematical models. Whilst still nascent, such an approach has proved useful in neuroscience in the form of biologically annotated connectomes [40], which are human brain networks obtained from structural imaging where each node (region) is imbued with additional information such as cell-type density, gene transcription, receptor density, and myelination. Open-source software, such as neuromaps [41], facilitates the integration of this multimodal data into richer structural representations. Such heterogeneity has been used to develop biophysical models of whole-brain dynamics that incorporate regional heterogeneity through variations in gene expression [42] or synaptic strength [43] for added biological complexity and closer agreement with observed neural dynamics.
@@
In an attempt to tackle the complexity of biological systems, the time has come to apply these frameworks more concretely to problems in biology. We are not arguing here that complex systems approaches to biology are entirely novel, but instead that the primary focus on emergent phenomena and the emphasis on universal phenomena have limited the accuracy and predictive power of these models. In turn, the mainstream within these fields may replace reductionist abstractions with richer representations that are capable of incorporating the heterogeneity seen in experiments. To achieve this goal, integration with multimodal data is crucial, but it is worth noting the scaling challenge that has proved limiting to this point. Given a system with $N$ elements, the number of pairwise interactions scales as $N^{2}$. If we introduce higher-order or multilayer interactions, the parameter space scales at an even faster rate $\sim N^{b}$. As a result, such over-parametrised models are plagued by both a lack of sufficient multimodal data and the inference algorithms needed to fit them. However, as we progress into the age of big data, modern experimental techniques are able to record multimodal data on an unprecedented scale, particularly in biological systems [44, 45]. Moreover, novel techniques in computational statistics and ML are able to perform parameter inference for such complex models using observed data [46, 47]. Unlike the purely statistical models typically found in ML, these systems have parameters that remain both physically constrained and meaningful. This ensures the interpretability of complex dynamical models and allays the usual concern about von Neumann’s elephant [^2] Whilst often attributed to von Neumann, the quote “…with four parameters I can fit an elephant, and with five I can make him wiggle his trunk.” first appears in an article by Freeman Dyson, where he quotes Enrico Fermi, who is in turn quoting von Neumann [48]. and his wiggly trunk [48]. With accurate calibration on experimental data and efficient simulation on HPCs, detailed, heterogeneous and elaborate models of complex systems can make the leap from generic and descriptive to mechanistic and predictive.
@@ II-A Integrating many time- and length-scales
The situation is completely different for biological systems that are best characterised by interacting spatiotemporal scales [49]. Focusing on a single scale, naturally limits our ability to explain certain phenomenon. For example, neurodegenerative diseases, like Alzheimer’s disease (AD), evolve due to the spread of toxic proteins over the course of many years [50]. However, the symptoms of these conditions manifest themselves in cognitive decline stemming from pathological neural activity, on the order of $10^{-3}-10^{0}$ seconds [51]. In order to understand the mechanisms by which AD causes cognitive decline, it is insufficient to only model either neural activity or toxic protein spread in isolation. Fig. 3 highlights the many processes underlying the development of AD along with their respective time- and length-scales. Recent studies have presented an approach that attempts to integrate these time-scales [52, 53]. First, they model network degradation due to toxic-protein spread, freezing the state of the network at distinct snapshots to then form the topology underlying a model of fast neural activity.
@@
Figure 3: Spatiotemporal scales in Alzheimer’s disease. Complex biological processes, such as the development of neurodegenerative diseases such as Alzheimer’s disease (AD), evolve over a range of interacting spatiotemporal scales. Focussing on AD, protein misfolding ultimately leads to the spread of toxic tau and amyloid-$\beta$ proteins through axonal fibres, which is partially cleared through cell regulation. Ultimately, this leads to cell death and pathological neural dynamics. This manifests in tau and amyloid pathology and, finally, cognitive decline which can be combatted with therapy. A $M$ of this process must take into account the full range of spatiotemporal scales in order to build a mechanistic understanding of AD.
@@ III-A Agent-based models
Mathematical models are often appreciated for their elegance and parsimony. In particular, such models can often be treated analytically leading to great insight. As models become more complex, they lose analytical tractability and numerical methods of analysis must be used. One framework for modelling complex systems that hinges on direct simulation, is the use of agent-based models (ABMs), which have been embraced in economics [54, 55], computational social science [56], and mathematical biology [57, 58].
The typical structure of ABMs is the time evolution of a micro-scale model that defines the characteristics of individual elements and their interactions. Examples of such systems include models of the spread of disease through populations and ecological dynamics or collective motion with agents being individuals or populations (Panel A, Fig. 4). Unlike traditional models, ABMs call for a different approach to mathematical analysis, one that is often realised through exhaustive simulation and data-driven calibration. For this reason, such models were previously found to be computationally restrictive. However, following advancements in scientific computing, large-scale simulations of ABMs have become both tractable and attractive [59].
The adoption of ABMs requires a shift in attitudes towards mathematical models. In particular, the proponents of such models argue that the complexity of many systems requires a shift towards a purely simulation-based science, where additional features and realism are favoured over parsimony and elegance [60]. The current inclination toward a minimal model is often justified via Occam’s razor, the philosophical argument that the simplest model is often the best. However, the lack of success of such minimal models in biology, compounded with the success of overparameterised ML models, has called into question this long-standing perspective [61, 62]. Whilst ML models are able to discover a multiplicity of latent features that have predictive power, a parsimonious mathematical modeller must instead identify the handful that may be important. Whilst not without critics [63], more complex models relieve the modeller of some of this choice by allowing them to include sufficient complexity at the level of model definition. Moreover, it is important to note the differences between an ABM and a ML model. Whilst both lack analytical tractability and rely on computationally-intensive processes, the latter may be considered to be a ‘black-box’, whose mechanisms remain somewhat opaque, whilst the former attempts to model the basic physical mechanisms governing a system. This poses additional advantages, such as the ability to test realistic interventions in a system due to the model’s interpretability [64].
@@ III-B Digital twins
Another simulation-based framework for the modelling of complex systems, is the digital twin (DT). A DT is an adaptive model that is designed to virtually emulate the behaviour of a physical system [67]. Typically such a model is implemented on a computer. DTs have been considerably useful in engineering where they can be used to simulate the behaviour of a system under a range of conditions. Successful examples include spacecraft during the Apollo missions [68], as well as other mechanical and aerospace engineering systems [69]. In some cases, a DT is composed of an ABM, such as in models of cities. Similarly to ABMs, whilst DTs have an underlying mathematical formulation, they require exhaustive simulation and HPC. Moreover, they are often characterised by more engineered interfaces such as virtual or augmented reality (VR/AR), for use by stakeholders other than the mathematical modeller.
@@
In biology, DTs have sparked particular interest in the field of precision medicine [70]. Precision medicine refers to a treatment paradigm that considers the individual characteristics and data of a patient. This can be realised through the use of a personalised DT that uses biomedical data to calibrate a simulation of a specific patient’s system to test interventions in-silico [^3] As opposed to experiments in-vitro (in glass i.e. in a laboratory) or in-vivo (in a living system), in-silico refers to a third paradigm where experiments can be run on a computer., as illustrated in Panel B, Fig. 4. Two particular systems showing great promise are the heart [71] and the brain [72]. In particular, The Virtual Brain (TVB), more specifically The Virtual Epileptic Patient (VEP), is a project leveraging detailed biophysical models and HPCs, that predicts personalised estimates of epileptogenic zone networks [73]. The empirical, actionable success of this approach stands in contrast with the dominant paradigm of complex systems neuroscience that focuses on simple oscillatory models of neural dynamics [74] and has yet to realise a concrete medical application [75]. Nevertheless, it would be naive to assume that the pure emulation of a system is either sufficient or necessary to understand it. The Human Brain Project [^4] Continuing now as the Blue Brain Project., was one such attempt to emulate the whole brain from individual neurons upwards [76], that, despite some notable results at the micro-scale [77], completely failed to deliver on its original promises [78]. On the other hand, whilst models like the VEP can have strong predictive power, this may come at the cost of interpretability. Important neuroscientific phenomena such as synchronisation or state-switching can be obscured by the more complex model whilst they would be obvious in an oscillator network or hidden Markov model. Moreover, they require large datasets of personalised neural recordings, therefore relying primarily on EEG, as opposed to more expensive, but much richer, modalities such as fMRI or MEG [^5] Electroencephalography (EEG), functional magnetic resonance imaging (fMRI) and magnetoencephalography (MEG) are methods of recording neural activity in humans.. Finally, the VEP requires large computational resources, as opposed to simplified models, which can be yield some interesting and important insights at the ‘laptop-level’.
At their best, DTs incorporate the necessary complexity to answer specific questions, test interventions, and simulate different environments, limited by the available data, rather than a perfectly faithful reconstruction of an entire system which achieves very little in scientific terms.